Calculate the number of atoms sampled in a 2x2x2,um-cubic in the surface of Nickel

1
Expert's answer
2022-03-02T03:11:08-0500

8 cubic micrometer

= 8× 10^-12cm3 or g

Number of moles = mass/ rfm

= 8× 10^-12/58.7= 1.36× 10^-13

Avagadro constant= 6.022× 10^23

6.022× 10^23 × 1.36×10^-13

= 8.2× 10^10 atoms